OpenStudy (anonymous):

an executive committee consists of 13 members: 5 men and 8 women. 3 members are selected at random to attend a meeting in Hawaii the names are drawn from a hat what is the probability that all 3 selected are women

OpenStudy (amistre64):

its a multiplication setup:\[\frac{w}{n}\frac{w-1}{n-1}\frac{w-2}{n-2}\]

OpenStudy (amistre64):

cant recall if theres a 13 c k tho

OpenStudy (blurbendy):

8/13 * 7/12 * 6/11

OpenStudy (sumi29):

Simple solution: 8c3/13c3 = 28/143. That's the final answer.

OpenStudy (amistre64):

i think this are Ps ... unless it simplifies

OpenStudy (amistre64):

lol, they are the same result

OpenStudy (amistre64):

\[\frac{nCr}{mCr}=\frac{(nPr)/r!}{(mPr)/r!}=\frac{nPr}{mPr}\] nice

OpenStudy (sumi29):

Yeah they are the same, I just wrote C coz it sprang up my mind. LOL. Anyways I hope the poster got the idea.